You have been asked to compare two alternatives, X and Y. MARR is 8%

EOY X Y

0

-4800 -4200
1 2000 2000
2 2500 2250
3 3000 2500
NPV @ 8% 1576.70

1565.40

  1. Which alternative would you recommend if using a Net Present Value measure of worth? Why?

b. Which alternative would you recommend if using a Benefit/Cost measure of worth? Why?
Answer #1

a. As NPV of X is more than Y, so X should be selected

b. We need to perform incremental B/C analysis

Incremental initial cost (X-Y) = 4800 - 4200 = 600

Incremental cash flow in EOY 1 (X-Y) = 2000 - 2000 = 0

Incremental cash flow in EOY 2 (X-Y) = 2500 - 2250 = 250

Incremental cash flow in EOY 3 (X-Y) = 3000 - 2500 = 500

Incremental B/C ratio (X-Y) = 250 * (P/G, 8%,3) / 600

= 250 * 2.4450 / 600

= 611.25 / 600

= 1.02

As Incremental B/C ratio (X-Y) is more than 1, X should be selected